OpenStudy (anonymous):

3 pens and 2 notebooks cost $8.25. 2 pens and 3 notebooks cost $8 How much would 2 pens and 2 notebooks cost?

OpenStudy (anonymous):

My answer for 1 Notebook is 1.5 1 pen is 1.75. But i'm sure that it's wrong.

OpenStudy (anonymous):

Pens can't cost more than notebooks... :P

OpenStudy (anonymous):

Well let's set up 2 equations: 3x+2y=8.25 and 2x+3y=8 where x is how much a pen costs and y is how much a notebook costs. now we can solve for one and plug in to the other 2x=8-3y x=4-1.5y Plug in to the other equation 3x+2y=8.25 3(4-1.5y)+2y=8.25 And just solve from there for y, and then you can plug the number you got for y back into either equation to get x :) hope this helps!

OpenStudy (anonymous):

thakns (: